Nine times the difference of a number and 22 is the same as 22 increased by nine times the number plus twice the number. Find the number.
Answer by addingup(3677)

Let the number be x:
9(x-22) = 22+9x+2x
9x-198 = 22+11x
-2x = 220
x = -110
>>>>>>>>>>>>>>>>
Check:
9(-110-22) = 22+9(-110)+2(-110)
9(-132) = 22+-990+-220
-1188 = -1188 Correct
:
John